# A physics-compliant diagonal representation for wireless channels parametrized by beyond-diagonal reconfigurable intelligent surfaces

### v1

@@ Abstract
The parametrization of wireless channels by so-called “beyond-diagonal reconfigurable intelligent surfaces” (BD-RIS) is mathematically characterized by a matrix whose off-diagonal entries are partially or fully populated. Physically, this corresponds to tunable coupling mechanisms between the RIS elements that originate from the RIS control circuit. Here, we derive a physics-compliant diagonal representation for BD-RIS-parametrized channels. Recognizing that the RIS control circuit, irrespective of its detailed architecture, can always be represented as a multi-port network with auxiliary ports terminated by tunable individual loads, we physics-compliantly express the BD-RIS-parametrized channel as a multi-port chain cascade of i) radio environment, ii) static parts of the control circuit, and iii) individually tunable loads. Thus, the cascade of the former two systems is terminated by a system that is mathematically always characterized by a diagonal matrix. This physics-compliant diagonal representation implies that existing algorithms for channel estimation and optimization for conventional (“diagonal”) RIS can be readily applied to BD-RIS scenarios. We demonstrate this in an experimentally grounded case study. Importantly, we highlight that, operationally, an ambiguous characterization of the cascade of radio environment and the static parts of the control circuit is required, but not the breakdown into the characteristics of its two constituent systems nor the lifting of the ambiguities. Nonetheless, we demonstrate how to derive or estimate the characteristics of the static parts of the control circuit for pedagogical purposes. The diagonal representation of BD-RIS-parametrized channels also enables their treatment with coupled-dipole-based models like PhysFad. In passing, we furthermore derive the assumptions under which the physics-compliant BD-RIS model for generic (potentially rich-scattering) radio environments simplifies to the widespread linear cascaded model.
@@ I Introduction
The properties of the channels used for wireless communications are traditionally imposed by the preexisting radio environment. With the advent of the “smart radio environment” [2, 3, 4, 5, 6], the wireless channels become to some extent controllable because they are deterministically parametrized by so-called reconfigurable intelligent surfaces (RISs) that are placed within the radio environment. An RIS is an array of elements with individually adjustable scattering properties. In practise, an RIS often takes the form of an array of patch antennas, each terminated by an individually tunable load impedance. Both in simplified cascaded channel models as well as in physics-compliant end-to-end channel models, the tunable RIS configuration appears along the diagonal of an otherwise empty matrix (see details in Sec. II below). For this reason, we refer to such conventional RIS designs as “diagonal RIS” (D-RIS) in this paper.
@@
In this paper, we introduce a thus far overlooked and at first sight counter-intuitive perspective on the representation of BD-RIS-parametrized wireless channels. Specifically, we propose a physics-compliant end-to-end representation of BD-RIS-parametrized wireless channels involving a diagonal tunable matrix. Our fundamental insight is that irrespective of the BD-RIS load circuit design, the latter is itself a multi-port network partially terminated by tunable individual loads. Hence, the BD-RIS-parametrized wireless channel is a multi-port chain cascade of three systems:
@@
    the static parts of the load circuit (from here on referred to as load circuit for conciseness), and
@@
    the individual tunable loads.
The conventional approach considers the radio environment whose auxiliary ports associated with the RIS elements are terminated by the cascade of load circuit and individual tunable loads that is characterized by a “beyond diagonal” matrix. Our alternative approach considers the cascade of radio environment with load circuit; the auxiliary ports of this cascade associated with tunable lumped elements within the load circuit are then terminated by individual tunable loads that are characterized by a diagonal matrix.
The most important implication of our formulation from a signal-processing perspective is that, in principle, our formulation obviates the need to develop new algorithms for BD-RIS, be it for end-to-end channel estimation or to optimize the BD-RIS configuration; instead, existing algorithms for D-RIS can be directly applied to BD-RIS by replacing the radio environment with the cascade of radio environment and load circuit. Note that for any practical realization of a BD-RIS, the characteristics of the load circuit will necessarily be well defined and hence known or characterizable.
@@ I-A Contributions
The contributions of the present paper are summarized as follows:
  1. 1.
@@
    Physics-compliant diagonal representation of BD-RIS-parametrized channels. We provide a physics-compliant end-to-end formulation in terms of scattering parameters that is more compact than the formulation in terms of impedance parameters that we proposed in the associated conference paper [1] because the wireless channel matrix is an off-diagonal block of a scattering matrix (see Sec. II). Moreover, the scattering-parameter formulation can be applied to canonical BD-RIS load circuit designs for which an impedance or admittance formulation would fail.
  2. 2.
    Identification of assumptions necessary to derive simplified cascaded model from physics-compliant model. We clarify the assumptions under which one can derive the widespread simplified cascaded channel model for BD-RIS from the fully physics-compliant end-to-end channel model based on multi-port network theory. Previously, only the assumptions under which one can derive the cascaded model for D-RIS based on PhysFad were elaborated in Ref. [17].
  3. 3.
    Identification of the load circuit scattering matrix $\mathbf{S}^{\mathrm{LC}}$. We analytically derive $\mathbf{S}^{\mathrm{LC}}$ for load circuits based on ideal T or $\pi$ networks with tunable impedances. We also explain how $\mathbf{S}^{\mathrm{LC}}$ can be obtained numerically in a single full-wave simulation given a concrete practical load circuit, as well as why it cannot be estimated experimentally without ambiguities and why that does not pose any problem.
  4. 4.
    Experimentally grounded demonstration of physics-compliant end-to-end estimation and optimization of BD-RIS-parametrized channels using D-RIS algorithms. We demonstrate in an experimentally grounded case study that D-RIS algorithms for channel estimation and optimization can be readily applied to BD-RIS-parametrized channels. We use scattering measurements obtained within a complex experimental environment and consider that the RIS elements are connected in groups of two to ideal $\pi$ networks whose impedances can take two possible values corresponding to those of a commercial PIN diode. To the best of our knowledge, this is the first experimentally grounded study on BD-RIS. We demonstrate that we can estimate the parameters of the proposed physics-compliant diagonal representation such that we can subsequently maximize the received signal strength indicator (RSSI) of a single-input single-output (SISO) link. We highlight ambiguities in the estimated parameters and demonstrate that, operationally, they do not matter. Importantly, we only characterize the cascade of radio environment and load circuit, without ever breaking it down into its two constituents. In other words, determining $\mathbf{S}^{\mathrm{LC}}$ is in fact not required (except for pedagogical reasons in the previous point).
  5. 5.
    Treatment of BD-RIS within coupled-dipole-based physics-compliant channel models like PhysFad [18, 19]. We clarify that the proposed physics-compliant diagonal representation also answers the open question of how coupled-dipole-based models can cope with the BD-RIS concept. The dipoles represent the primary entities, which are the transmitter and receiver ports as well as the tunable lumped elements. The dipoles are coupled to each other via background Green’s functions lumping together complex scattering in the environment (including structural scattering) in the D-RIS case and, capturing the scattering within the cascade of the radio environment and the load circuit in the BD-RIS case.
@@ I-B Outline
The remainder of this paper is organized as follows. In Sec. II, we develop the physics-compliant multi-port network chain cascade description of BD-RIS-parametrized wireless channels, free of any approximations, that yields the diagonal representation at the core of this paper. In Sec. III, we identify the scattering properties of the load circuit analytically for two canonical ideal load circuits, and we discuss how to determine it numerically or experimentally for practical load circuits. In Sec. IV, we demonstrate the direct applicability of existing channel estimation and optimization algorithms for D-RIS in an experimentally grounded case study of a BD-RIS-parametrized rich-scattering SISO link. In Sec. VI, we briefly conclude.
@@ II Theory
The generic multi-port chain cascade formulation for BD-RIS-parametrized wireless channels developed in this section is based on multi-port network theory which can equivalently be expressed in terms of scattering, impedance or admittance parameters. We use scattering parameters in this section for two reasons: first, scattering parameters yield a more compact formulation since the wireless channels are scattering coefficients; second, scattering parameters are suitable to describe canonical examples of BD-RIS load circuits that constitute pathological special cases in which impedance parameters cannot be defined. The general impedance formulation can be found in our associated conference paper [1]. A discussion about the relation to equivalent coupled-dipole-based physics-compliant channel formulations like PhysFad [18, 20, 19, 17] is provided in Sec. V.
@@
II-A Problem statement
@@
Building on a decade-old circuit theory of communication [22, 23], the literature contains by now many proposed formulations for physics-inspired models of D-RIS-parametrized channels based on multi-port network theory [7, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36], as well as three formulations for BD-RIS-parametrized channels [37, 1, 16]. These numerous formulations differ regarding the approximations and assumptions that they make, especially regarding the nature of the radio environment. In the present paper, we do not make any assumptions beyond those stated in the previous paragraph, ensuring that our formulation is fully physics-compliant. Moreover, we work with the most compact and universal multi-port description of an arbitrarily complex smart radio environment that describes the latter with the scattering matrix $\mathbf{S}^{\mathrm{RE}}\in\mathbb{C}^{N_{\mathrm{RE}}\times N_{\mathrm{RE}}}$ that lumps together all scattering effects, including scattering objects and walls (irrespective of whether they are point-like or not) and structural scattering [38, 39, 40] by antennas and RIS elements. Importantly, the number of parameters is hence independent of the radio environment’s complexity and only depends on the number of attached ports $N_{\mathrm{RE}}=N_{\mathrm{T}}+N_{\mathrm{R}}+N_{\mathrm{S}}$, where $N_{\mathrm{T}}$, $N_{\mathrm{R}}$ and $N_{\mathrm{S}}$ denote the number of transmitting antennas, receiving antennas and RIS elements, respectively. Moreover, as long as $\mathbf{S}^{\mathrm{RE}}$ is known or can be estimated [^2] See Ref. [19, 30, 16] for various experimentally realized methods for end-to-end channel estimation, with varying degrees of parameter ambiguities. Irrespective of these ambiguities, the end-to-end wireless channels are correctly predicted as a function of D-RIS configuration in all cases. Further discussion is provided in Sec. IV-B., no explicit description of the radio environment (in terms of geometry, materials, etc.) is necessary. These two important aspects first emerged in PhysFad-based works [20, 19] and were subsequently transposed to an equivalent multi-port network formulation in Ref. [28] where all required model parameters were extracted from a single full-wave simulation for an overwhelmingly complicated D-RIS-parametrized on-chip radio environment featuring reflective walls, extended dielectric layers and strong structural scattering.
[…]
Overall, as illustrated in Fig. 1, the BD-RIS-parametrized radio environment is hence a chain cascade of radio environment (RE), load circuit (LC), and individual tunable loads (IL). To work out the end-to-end wireless channel matrix $\mathbf{H}\in\mathbb{C}^{N_{\mathrm{R}}\times N_{\mathrm{T}}}$, we can proceed in two distinct manners. As shown in Fig. 1, the conventional perspective (albeit never explicited in this manner so far) first evaluates the cascade of LC and IL, yielding a generally “beyond-diagonal” scattering matrix that terminates the radio environment’s auxiliary ports associated with the RIS elements. The alternative perspective, highlighted and leveraged in the present paper, first cascades RE and LC, which yields a multi-port network whose auxiliary ports associated with the tunable lumped elements are then terminated by IL which is characterized by a scattering matrix that is always diagonal. We elaborate on these two perspectives in the subsequent subsections.
[…]
II-B Conventional perspective
As mentioned, for the conventional perspective, we begin by evaluating the cascade of LC and IL that is characterized by the scattering matrix $\mathbf{S}^{\mathrm{L}}\in\mathbb{C}^{N_{\mathrm{S}}\times N_{\mathrm{S}}}$ which terminates the RIS element auxiliary ports of the radio environment. Let $\mathbf{S}^{\mathrm{IL}}\in\mathbb{C}^{N_{\mathrm{C}}\times N_{\mathrm{C}}}$ be a diagonal matrix containing the $N_{\mathrm{C}}$ reflection coefficients of the individual tunable loads. We denote by $\bar{\mathcal{S}}$ and $\mathcal{C}$ the sets of port indices of the load circuit associated with RIS elements and tunable individual loads, respectively. Then,
$\mathbf{S}^{\mathrm{L}}=\mathbf{S}^{\mathrm{LC}}_{\bar{\mathcal{S}}\bar{\mathcal{S}}}+\mathbf{S}^{\mathrm{LC}}_{\bar{\mathcal{S}}\mathcal{C}}\left(\left(\mathbf{S}^{\mathrm{IL}}\right)^{-1}-\mathbf{S}^{\mathrm{LC}}_{\mathcal{CC}}\right)^{-1}\mathbf{S}^{\mathrm{LC}}_{\mathcal{C}\bar{\mathcal{S}}},$ (1)
[…]
Given $\mathbf{S}^{\mathrm{L}}$, we can then determine the scattering matrix $\tilde{\mathbf{S}}\in\mathrm{C}^{N_{\mathrm{A}}\times N_{\mathrm{A}}}$, where $N_{\mathrm{A}}=N_{\mathrm{T}}+N_{\mathrm{R}}$, that is “measurable” via the antennas in the radio environment:
@@ II-B Conventional perspective
where $\mathcal{A}$ and $\mathcal{S}$ denote the sets of port indices of the radio environment associated with the antennas and RIS elements, respectively. The wireless channel matrix $\mathbf{H}$ is then simply an off-diagonal block of $\tilde{\mathbf{S}}$:
$\mathbf{H}=\left[\tilde{\mathbf{S}}\right]_{\mathcal{RT}},$ (3)
where $\mathcal{R}$ and $\mathcal{T}$ denote the sets containing the port indices associated with receiving and transmitting antennas, respectively. For completeness, we note that $\mathcal{A}=\mathcal{T}\cup\mathcal{R}$.
Following the procedure detailed for D-RIS based on PhysFad in Ref. [17], we can rewrite the matrix inversion in Eq. (2) as infinite sum of matrix powers:
$\begin{split}\tilde{\mathbf{S}}&=\mathbf{S}^{\mathrm{RE}}_{\mathcal{AA}}+\mathbf{S}^{\mathrm{RE}}_{\mathcal{AS}}\left(\left(\mathbf{S}^{\mathrm{L}}\right)^{-1}-\mathbf{S}^{\mathrm{RE}}_{\mathcal{SS}}\right)^{-1}\mathbf{S}^{\mathrm{RE}}_{\mathcal{SA}}\\ &=\mathbf{S}^{\mathrm{RE}}_{\mathcal{AA}}+\mathbf{S}^{\mathrm{RE}}_{\mathcal{AS}}\left(\mathbf{I}_{N_{\mathrm{S}}}-\mathbf{S}^{\mathrm{L}}\mathbf{S}^{\mathrm{RE}}_{\mathcal{SS}}\right)^{-1}\mathbf{S}^{\mathrm{L}}\mathbf{S}^{\mathrm{RE}}_{\mathcal{SA}}\\ &=\mathbf{S}^{\mathrm{RE}}_{\mathcal{AA}}+\mathbf{S}^{\mathrm{RE}}_{\mathcal{AS}}\left(\sum_{k=0}^{\infty}\left(\mathbf{S}^{\mathrm{L}}\mathbf{S}^{\mathrm{RE}}_{\mathcal{SS}}\right)^{k}\right)\mathbf{S}^{\mathrm{L}}\mathbf{S}^{\mathrm{RE}}_{\mathcal{SA}},\end{split}$ (4)
where the convergence of the infinite series is guaranteed by the fact that we work with passive systems. Specifically, passivity ensures that the spectral radii of $\mathbf{S}^{\mathrm{L}}$ and $\mathbf{S}^{\mathrm{RE}}_{\mathcal{SS}}$ do not exceed unity, and hence, due to submultiplicativity, the spectral radius of $\mathbf{S}^{\mathrm{L}}\mathbf{S}^{\mathrm{RE}}_{\mathcal{SS}}$ cannot exceed unity. Upon truncation of the infinite series in the last line of Eq. (4) after its first term, i.e., $k=0$, we obtain
$\tilde{\mathbf{S}}\approx\mathbf{S}^{\mathrm{RE}}_{\mathcal{AA}}+\mathbf{S}^{\mathrm{RE}}_{\mathcal{AS}}\mathbf{S}^{\mathrm{L}}\mathbf{S}^{\mathrm{RE}}_{\mathcal{SA}}.$ (5)
Upon inserting Eq. (5) into Eq. (3), we get

### v2

@@ Abstract
The parametrization of wireless channels by so-called “beyond-diagonal reconfigurable intelligent surfaces” (BD-RIS) is mathematically characterized by a matrix whose off-diagonal entries are partially or fully populated. Physically, this corresponds to tunable coupling mechanisms between the RIS elements that originate from the RIS control circuit. Here, we derive a physics-compliant diagonal representation for BD-RIS-parametrized channels. We recognize that any RIS control circuit can always be separated into its static parts (SLC) and a set of tunable individual loads (IL). Therefore, a BD-RIS-parametrized channel results from the chain cascade of three systems: i) radio environment (RE), ii) SLC, and iii) IL. RE and SLC are static non-diagonal systems whose cascade K is terminated by the tunable diagonal system IL. This physics-compliant representation in terms of K and IL is directly analogous to that for conventional (“diagonal”) RIS (D-RIS). Therefore, scenarios with BD-RIS can also readily be captured by the physics-compliant coupled-dipole model PhysFad, as we show. In addition, physics-compliant algorithms for system-level optimization with D-RIS can be directly applied to scenarios with BD-RIS. We demonstrate this important implication of our conceptual finding in a case study on end-to-end channel estimation and optimization in a BD-RIS-parametrized rich-scattering environment. Our case study is the first experimentally grounded system-level optimization for BD-RIS: We obtain the characteristics of RE and IL from experimental measurements and a commercial PIN diode, respectively. Altogether, our physics-compliant diagonal representation for BD-RIS enables a paradigm shift in how practitioners in wireless communications and signal processing implement system-level optimizations for BD-RIS because it enables them to directly apply existing physics-compliant D-RIS algorithms.
@@ I Introduction
The properties of the channels used for wireless communications are traditionally imposed by the preexisting radio environment. With the advent of the “smart radio environment” [2, 3, 4, 5, 6], the wireless channels become to some extent controllable because they are deterministically parametrized by so-called reconfigurable intelligent surfaces (RISs) that are placed within the radio environment. An RIS is an array of elements with individually adjustable scattering properties. In practice, an RIS often takes the form of an array of patch antennas whose ports are terminated by individual tunable lumped elements. Both in simplified cascaded channel models as well as in physics-compliant end-to-end channel models, the tunable RIS configuration appears along the diagonal of an otherwise empty matrix (see details in Sec. II below). For this reason, we refer to such conventional RIS designs as “diagonal RIS” (D-RIS) in this paper.
@@
To avoid confusion about the terminology, we emphasize that we use the terms “D-RIS” and “BD-RIS” to characterize the load circuit of the RIS without making any statement about whether the end-to-end channel model neglects or accounts for mutual coupling between RIS elements arising due to their spatial proximity or reverbreation within the radio environment. In other words, a “D-RIS” load circuit architecture does not imply that a simplified cascaded model neglecting mutual coupling is assumed.
@@
Most research efforts on D-RIS and BD-RIS are primarily focused on system-level optimizations, assuming simplified cascaded channel models. Because the tunability of BD-RIS is represented not by a diagonal but a more densely (potentially fully populated) matrix, the mathematical structure of the channel parametrization appears to differ from that in the D-RIS case, motivating recent efforts to conceive special algorithms for system-level optimization dedicated to BD-RIS architectures. Meanwhile, more and more researchers seek to model and optimize RIS-assisted wireless communications in a way that is electromagnetically consistent, by accounting for various mutual coupling effects. We summarize below these efforts based on multi-port network theory (Sec. II-A) or coupled-dipole formalisms (Sec. II-C1).
In this paper, we introduce a thus far overlooked and at first sight counter-intuitive perspective on the representation of BD-RIS-parametrized wireless channels. Specifically, we propose a physics-compliant end-to-end representation of BD-RIS-parametrized wireless channels involving a diagonal tunable matrix. Our fundamental insight is that irrespective of the BD-RIS load circuit design, the latter can be separated into a static part and a tunable part that typically relies on tunable lumped elements. Therefore, the BD-RIS-parametrized wireless channel is a chain cascade of three systems: [^2] Mathematically, the physics-compliant evaluation of a chain cascade of multiple systems is generally more complicated than a chain multiplication of three matrices – see Sec. II.
@@
    SLC: the static parts of the load circuit, and
@@
    IL: the individual tunable loads.
Only IL is tunable, while RE and SLC are static. Moreover, only IL is characterized by a diagonal scattering matrix, while RE and SLC are generally characterized by more densely populated non-diagonal scattering matrices. Therefore, we propose to group together the two static and non-diagonal systems RE and SLC, singling out the tunable diagonal system IL. Thereby, we obtain a representation of the end-to-end wireless channel involving a static non-diagonal system (cascade of RE and SLC) and a tunable diagonal system (IL). This representation is analogous to the case of D-RIS in which the static non-diagonal system is simply RE and the tunable diagonal system is IL [^3] While the number of tunable lumped elements equals the number of RIS elements in a D-RIS architecture, this is generally not the case in a BD-RIS architecture – see Sec. II.. In contrast to our proposal, the conventional representation of BD-RIS-parametrized channels groups together the static non-diagonal system SLC and tunable diagonal system IL, yielding a representation of the end-to-end channel involving a static non-diagonal system (RE) and a tunable non-diagonal system (cascade of SLC and IL).
The most important implication of our formulation from the perspectives of wireless communications and signal processing is that, in principle, our formulation obviates the need to develop new algorithms specifically dedicated to BD-RIS for system-level optimizations. Instead, existing physics-compliant algorithms for D-RIS can be directly applied to BD-RIS by replacing RE in the D-RIS case with the cascade of RE and SLC in the BD-RIS case.
@@ I-A Contributions
Our main conceptual (theoretical) contribution is the introduction of a physics-compliant diagonal representation of BD-RIS-parametrized channels. We recognize that the BD-RIS load circuit can be split into SLC and IL, implying that a BD-RIS-parametrized channel arises from the chain cascade of three systems: RE, SLC, IL. Hence, the cascade of RE and SLC, named K, is terminated by the tunable diagonal system IL – directly analogous to physics-compliant models of D-RIS-parametrized channels. More specific details regarding our conceptual contribution are as follows:
@@
  • In contrast to the associated conference paper [1] based on impedance parameters, we use scattering parameters. Thereby, our formulation is more compact and applies even to canonical BD-RIS load circuit designs for which an impedance or admittance formulation would fail.
  • We establish the equivalence between the formulation in terms of scattering parameters and the coupled-dipole-based PhysFad model for D-RIS [17]. Thereby, we clarify how BD-RIS can be treated within PhysFad.
  • We explain how the scattering parameters of SLC can be identified in various scenarios that can arise in theoretical, numerical and experimental studies. These scenarios range from canonical load circuits based on ideal T or $\pi$ networks with tunable impedances, via full-wave simulations, to experimental estimations.
  • We thoroughly explain under what assumptions the widespread cascaded channel model can be derived from the physics-compliant end-to-end model formulated in terms of scattering parameters for a channel parametrized by a D-RIS or BD-RIS.
Our main algorithmic contribution is the experimentally grounded demonstration of physics-compliant end-to-end estimation and optimization of a BD-RIS-parametrized channel using D-RIS algorithms. We evidence the most important implication of our conceptual finding in the realm of system-level optimization: Existing physics-compliant D-RIS algorithms can be directly applied to BD-RIS. More specific details regarding this contribution are as follows:
  • We perform end-to-end channel estimation in a rich-scattering BD-RIS-parametrized radio environment using an existing physics-compliant D-RIS algorithm. We highlight the existence and operational irrelevance of parameter ambiguities.
  • We perform RSSI maximization based on the estimated physics-compliant diagonal end-to-end channel model, achieving optimal performance as confirmed by an exhaustive brute-force search.
  • This is the first experimentally grounded study on system-level optimization for BD-RIS (all existing papers are purely theoretical). The assumed characteristics of RE are based on experimental measurements in a rich-scattering environment. The assumed characteristics of IL are based on a commercial PIN diode.
@@ I-B Outline
The remainder of this paper is organized as follows. In Sec. II, we develop the theory of our physics-compliant diagonal representation for BD-RIS: We start with background on multi-port network models for RIS-parametrized channels (Sec. II-A), then we present our key conceptual insight (Sec. II-B), and finally we transpose it to the PhysFad model (Sec. II-C). In Sec. III, we explain how to identify the scattering properties of SLC analytically for two canonical ideal load circuits, and we discuss how to determine it numerically or experimentally for practical load circuits. In Sec. IV, we demonstrate the direct applicability of existing physics-compliant D-RIS algorithms for end-to-end channel estimation and optimization in an experimentally grounded case study of a BD-RIS-parametrized rich-scattering environment. In Sec. V, we briefly conclude.
@@ I-C Notation
  • $\mathbf{1}_{m}$ denotes an $m$-element column vector whose entries are all unity.
@@
  • $\mathrm{vec}(\mathbf{A})$ denotes the vectorization of the matrix $\mathbf{A}$.
  • $\mathbf{S}^{\mathrm{X}}$ is the scattering matrix of the multi-port system X.
  • $\mathbf{S}_{\mathcal{FG}}^{\mathrm{X}}$ denotes the block of $\mathbf{S}^{\mathrm{X}}$ identified by the sets of row [column] indices $\mathcal{F}$ [$\mathcal{G}$].
@@ II Theory
The generic multi-port chain cascade formulation for BD-RIS-parametrized wireless channels developed in this paper is based on multi-port network theory which can equivalently be expressed in terms of scattering, impedance or admittance parameters (background is provided in Sec. II-A). Moreover, an equivalent coupled-dipole formalism coined PhysFad [17, 18, 19, 20] exists (background is provided in Sec. II-C1). We use scattering parameters as opposed to impedance or admittance parameters for two reasons: first, scattering parameters yield a more compact formulation since the wireless channels are scattering coefficients; second, scattering parameters are suitable to describe canonical examples of BD-RIS load circuits that constitute pathological special cases in which impedance parameters cannot be defined. The general impedance formulation can be found in our associated conference paper [1]. A discussion about the relation to PhysFad is provided in Sec. II-C.
@@
II-A Background on Multi-Port Network Models for RIS-Parametrized Channels
@@
Building on a decade-old circuit theory of communication [22, 23], the literature contains by now many proposed formulations for physics-inspired models of D-RIS-parametrized channels based on multi-port network theory [7, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36], as well as three formulations for BD-RIS-parametrized channels [37, 1, 16]. These numerous formulations differ regarding the approximations and assumptions that they make, especially regarding the nature of the radio environment. In the present paper, we do not make any assumptions beyond those stated in the previous paragraph, ensuring that our formulation is fully physics-compliant. Moreover, we work with the most compact and universal multi-port description of an arbitrarily complex smart radio environment that describes the latter with the scattering matrix $\mathbf{S}^{\mathrm{RE}}\in\mathbb{C}^{N_{\mathrm{RE}}\times N_{\mathrm{RE}}}$ that lumps together all scattering effects, including scattering objects and walls (irrespective of whether they are point-like or not) and structural scattering [38, 39, 40] by antennas and RIS elements. Importantly, the number of parameters is hence independent of the RE’s complexity and only depends on the number of attached ports $N_{\mathrm{RE}}=N_{\mathrm{T}}+N_{\mathrm{R}}+N_{\mathrm{S}}$, where $N_{\mathrm{T}}$, $N_{\mathrm{R}}$ and $N_{\mathrm{S}}$ denote the number of transmitting antennas, receiving antennas and RIS elements, respectively. Moreover, as long as $\mathbf{S}^{\mathrm{RE}}$ is known or can be estimated [^4] See Ref. [19, 30, 16] for various experimentally realized methods for physics-compliant end-to-end channel estimation, with varying degrees of parameter ambiguities. Irrespective of these ambiguities, the end-to-end wireless channels are correctly predicted as a function of D-RIS configuration in all cases. Further discussion is provided in Sec. IV-B., no explicit description of the RE (in terms of geometry, materials, etc.) is necessary. These two important aspects first emerged in PhysFad-based works [18, 19] and were subsequently transposed to an equivalent multi-port network formulation in Ref. [28] where all required model parameters were extracted from a single full-wave simulation for an overwhelmingly complicated D-RIS-parametrized on-chip RE featuring reflective walls, extended dielectric layers and strong structural scattering.
The RE’s $N_{\mathrm{S}}$ auxiliary ports associated with the $N_{\mathrm{S}}$ RIS elements are terminated by a tunable load circuit. Hence, the load circuit (denoted by L in the following) has $N_{\mathrm{S}}$ ports and is characterized by a scattering matrix $\mathbf{S}^{\mathrm{L}}\in\mathbb{C}^{N_{\mathrm{S}}\times N_{\mathrm{S}}}$. In the case of a D-RIS, $\mathbf{S}^{\mathrm{L}}$ is diagonal and contains the reflection coefficients of the $N_{\mathrm{S}}$ tunable lumped elements. In the case of a BD-RIS, $\mathbf{S}^{\mathrm{L}}$ is “beyond-diagonal” and includes the scattering matrix associated with the reconfigurable impedance network.
Given $\mathbf{S}^{\mathrm{RE}}$ and $\mathbf{S}^{\mathrm{L}}$, the standard “cascade loading” formula from multi-port network theory [41, 42, 43, 44] (see Appendix B.3 in Ref. [44] for a recent derivation) directly yields the resulting scattering matrix $\tilde{\mathbf{S}}\in\mathbb{C}^{N_{\mathrm{A}}\times N_{\mathrm{A}}}$, where $N_{\mathrm{A}}=N_{\mathrm{T}}+N_{\mathrm{R}}$, that is “measurable” via the antenna ports in the RE:
@@
where $\mathcal{A}$ and $\mathcal{S}$ denote the sets of port indices of the RE associated with the antennas and RIS elements, respectively.
The end-to-end wireless channel matrix $\mathbf{H}\in\mathbb{C}^{N_{\mathrm{R}}\times N_{\mathrm{T}}}$ is then simply an off-diagonal block of $\tilde{\mathbf{S}}$:
$\mathbf{H}=\left[\tilde{\mathbf{S}}\right]_{\mathcal{RT}}=\mathbf{S}^{\mathrm{RE}}_{\mathcal{RT}}+\mathbf{S}^{\mathrm{RE}}_{\mathcal{RS}}\left(\left(\mathbf{S}^{\mathrm{L}}\right)^{-1}-\mathbf{S}^{\mathrm{RE}}_{\mathcal{SS}}\right)^{-1}\mathbf{S}^{\mathrm{RE}}_{\mathcal{ST}},$ (2)
where $\mathcal{R}$ and $\mathcal{T}$ denote the sets of port indices associated with receiving and transmitting antennas, respectively, and $\mathcal{A}=\mathcal{T}\cup\mathcal{R}$. Importantly, we remark that in this physics-compliant end-to-end channel model, the dependence of $\mathbf{H}$ on the configuration of the RIS (encoded in $\mathbf{S}^{\mathrm{L}}$) is non-linear due to the matrix inversion that captures mutual coupling – irrespective of whether we consider a D-RIS or a BD-RIS.
To connect this physics-compliant model to the widespread simplified cascaded model, we rewrite the matrix inversion in Eq. (2) as infinite sum of matrix powers:
$\begin{split}\mathbf{H}&=\mathbf{S}^{\mathrm{RE}}_{\mathcal{RT}}+\mathbf{S}^{\mathrm{RE}}_{\mathcal{RS}}\left(\left(\mathbf{S}^{\mathrm{L}}\right)^{-1}-\mathbf{S}^{\mathrm{RE}}_{\mathcal{SS}}\right)^{-1}\mathbf{S}^{\mathrm{RE}}_{\mathcal{ST}}\\ &=\mathbf{S}^{\mathrm{RE}}_{\mathcal{RT}}+\mathbf{S}^{\mathrm{RE}}_{\mathcal{RS}}\left(\mathbf{I}_{N_{\mathrm{S}}}-\mathbf{S}^{\mathrm{L}}\mathbf{S}^{\mathrm{RE}}_{\mathcal{SS}}\right)^{-1}\mathbf{S}^{\mathrm{L}}\mathbf{S}^{\mathrm{RE}}_{\mathcal{ST}}\\ &=\mathbf{S}^{\mathrm{RE}}_{\mathcal{T}}+\mathbf{S}^{\mathrm{RE}}_{\mathcal{RS}}\left(\sum_{k=0}^{\infty}\left(\mathbf{S}^{\mathrm{L}}\mathbf{S}^{\mathrm{RE}}_{\mathcal{SS}}\right)^{k}\right)\mathbf{S}^{\mathrm{L}}\mathbf{S}^{\mathrm{RE}}_{\mathcal{ST}},\end{split}$ (3)
where the invertibility of $\left(\mathbf{I}_{N_{\mathrm{S}}}-\mathbf{S}^{\mathrm{L}}\mathbf{S}^{\mathrm{RE}}_{\mathcal{SS}}\right)$ and the convergence of the infinite series are guaranteed by two basic assumptions: First, RE is a passive sub-unitary system, i.e., it does not involve any gain but inevitably features some absorption and/or leakage, [^5] Absorption is the irreversible conversion of wave energy to other forms of energy like heat; gain would be the opposite. Meanwhile, leakage refers to wave energy exiting the system without passing through the considered ports. such that $\mathbf{S}^{\mathrm{RE}}_{\mathcal{SS}}$ is strictly sub-unitary; second, L is a passive system that may be considered lossless and hence unitary in some theoretical studies but would be sub-unitary in any practical realization due to inevitable absorption, such that $\mathbf{S}^{\mathrm{L}}_{\mathcal{SS}}$ is unitary or sub-unitary. Hence, due to submultiplicativity, the spectral radius of $\mathbf{S}^{\mathrm{L}}\mathbf{S}^{\mathrm{RE}}_{\mathcal{SS}}$ is always below unity and the series always converges. For the same reasons, $\mathbf{S}^{\mathrm{L}}\mathbf{S}^{\mathrm{RE}}_{\mathcal{SS}}$ can never have an eigenvalue of 1 such that $\left(\mathbf{I}_{N_{\mathrm{S}}}-\mathbf{S}^{\mathrm{L}}\mathbf{S}^{\mathrm{RE}}_{\mathcal{SS}}\right)$ can never be a singular matrix.
Upon truncation of the infinite series in the last line of Eq. (3) after its first term, i.e., $k=0$, we obtain